OpenStudy (anonymous):

check my answer? -3x + 4 <= x + 8 I got x<=1

ganeshie8 (ganeshie8):

looks you forgot to flip inequality when you divided it wid -4

ganeshie8 (ganeshie8):

lets do this step by step and see ok :)

OpenStudy (anonymous):

so you have to flip is even if you don't multiply or divide by 2 negatives? @ganeshie8

ganeshie8 (ganeshie8):

exactly, when u divide or multiply a negative, you need to flip the inequality

ganeshie8 (ganeshie8):

-3x + 4 <= x + 8

ganeshie8 (ganeshie8):

wats the first step you did ?

OpenStudy (anonymous):

oh darn i have a lot of problems to re-due now haha and i subtracted x from both sides @ganeshie8

ganeshie8 (ganeshie8):

oly that thing we need to remember when multiplying/dividing negative...

ganeshie8 (ganeshie8):

-3x + 4 <= x + 8 subtract x -3x + 4 <= x + 8 -x -x -4x + 4 <= 8

ganeshie8 (ganeshie8):

wats the next step ?

OpenStudy (anonymous):

sub 4 from both sides?

ganeshie8 (ganeshie8):

-4x + 4 <= 8 subtract 4 -4x + 4 <= 8 - 4 -4 -4x <= 4

ganeshie8 (ganeshie8):

next step ?

OpenStudy (anonymous):

divide both sides by -4

ganeshie8 (ganeshie8):

so we pause a sec. since we dividing both sides by negative, we need to flop the inequality

ganeshie8 (ganeshie8):

-4x <= 4 divide -4 -4x <= 4 /-4 /-4 x >= -1

OpenStudy (anonymous):

how would you graph that?

ganeshie8 (ganeshie8):

on a number line

ganeshie8 (ganeshie8):

goto -1, and draw an arrow to right side

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

thank you (: will you help me with another?

ganeshie8 (ganeshie8):